# Diversity of low-mass planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v2

@@ 5 Conclusions
For Earth-sized rocky planets like TRAPPIST-1e, the atmosphere during the magma ocean stage is likely CO-rich due to two key factors. First, CO(g) is thermodynamically favoured at 1800 K across a broad range of oxygen fugacities. Second, carbon-bearing species are significantly less soluble than hydrogen-bearing species—particularly H2O(g)—leading to an elevated atmospheric C/H, often exceeding the bulk C/H by several orders of magnitude. As a result, hydrogen-bearing species are depleted in the atmosphere, even when the mantle is only partially molten. In a fully molten mantle, the further depletion of hydrogen by the sequestration of H2O(g) into the liquid leads to atmospheres that are defined by entirely continuous mixing ratios of their constituent species, favouring atmospheres dominated by 1) heavy carbon species (predominantly CO(g)), 2) CO2(g) SO2(g)-bearing varieties at high O2s (IW=+3 to +4) or 3) H2(g) CH4(g)-bearing types at low O2 (below IW=-3).
@@
Based on the equilibrium results of atmospheres with molten mantles, we cool the atmospheres to 280 K and permit (if thermodynamically stable) the formation of water (oceans), graphite, -sulfur, and ammonium chloride. Unlike at high temperatures, the presence of condensates buffers gas mixing ratios to produce discrete ‘families’ of atmospheres. Notably, CO2-rich atmospheres are not only widespread (83% of cases) but also uniquely capable of supporting a wider variety of surface condensates compared to those dominated by other species. A small fraction (5%) are CH4-rich atmospheres that can sustain both water and graphite while roughly 8% are CO2-SO2-rich in equilibrium with native sulfur. The remainder (1%) are H2-rich. Depending on the conditions during early equilibration with a magma ocean, atmospheres rich in N2, HCl, or H2S can also host diverse surface condensates; however, such cases comprise only 1% of all cases. In short, water condensation buffers the partial pressures of hydrogen-bearing species in the atmosphere, such that even low C/H does not produce substantial hydrogen atmospheres except at the most reduced conditions where H2 can dominate. Thus, for both high and low temperature atmospheres, carbon always wins.
@@
Our simulations can be considered as geological reference cases—the expected state of a rocky planet after its earliest evolution. From this state, subsequent processes such as geological outgassing and atmospheric (photo)chemistry, dynamics, and escape can modify the atmosphere over geological timescales. Hence our results can be used to inform the priors and initial conditions of other models, as well as provide an observational baseline for rocky planets to determine if they have evolved away from their birth state. In this regard, the simulations can inform observational studies and instrument requirements for investigating the atmospheres of temperate rocky planets.
@@
For sub-Neptune-sized planets, we performed non-ideal gas-phase chemistry simulations including dominant C- and H-bearing species at conditions relevant to the magma ocean-envelope interface. The combination of solubility and non-ideality conspire to stabilise CH4 as a major gas species for a range of parameter choices that are reasonable given current observational constraints for sub-Neptunes. As for Earth-sized rocky planets, this result is a natural outcome of the high solubility of hydrogen-bearing species (notably molecular hydrogen for sub-Neptunes) compared to carbon-bearing species, a result that is exacerbated due to non-ideality at the high temperature (3000 K) and pressure (several GPa) conditions assumed to occur at the interface. Furthermore, the transition between a H2 and CH4 atmosphere can occur within the range of observational uncertainties in mass and radius, even for a fixed volatile budget and O2, thereby providing a new challenge to infer atmospheric composition. Given detections of carbon-bearing species in the atmospheres of sub-Neptunes, such as K2-18b and TOI-270d, our findings emphasise the need to include and realistically model gas envelopes throughout their extent.
@@
To achieve these results, we devised an open source code Atmodeller, built on JAX, to probe the nature of atmospheres above rocky planets from Earth size to sub-Neptune size. The code includes solubility laws, real gas equations of state, and thermodynamic data. Leveraging the functionality of the JAX Python library allows the code to be both user friendly and highly performant, and community development of the code is encouraged.
Acknowledgements
@@
We thank Daniel Kitzmann, Patrick Kidger, and Johanna Haffner for discussions. DJB, MAT, and PAS were supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract No. MB22.00033, a SERI-funded ERC Starting grant “2ATMO”. PAS also thanks the Swiss National Science Foundation (SNSF) through an Eccellenza Professorship (#203668). MAT is supported by NASA through the NASA Hubble Fellowship grant #HST-HF2-51545 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. KH acknowledges the FED-tWIN research program STELLA (Prf-2021-022) funded by the Belgian Science Policy Office (BELSPO) and the FWO research grant G014425N. MT acknowledges financial support from the Chair of Theoretical Astrophysics of Extrasolar Planets at LMU Munich.
@@ Appendix A Sampling strategy and parameter choices
The logarithmic sampling strategy adopted to explore the atmospheric diversity of TRAPPIST-1e is physically motivated as follows. Oxygen fugacity (O2) exerts a fundamental control on atmospheric composition, particularly through its influence on redox speciation, for example, determining the relative abundances of reduced versus oxidised species such as H2 and H2O. Thermodynamically, O2 is defined by an equilibrium reaction, and when referenced to Fe–FeO (iron-wüstite, IW) equilibrium (Equation E1) is conventionally expressed as a offset relative to the IW buffer. This convention naturally supports the use of a log-uniform prior, especially within a thermochemical framework where partial pressures, fugacities, and equilibrium constants combine multiplicatively (and thus additively in log-space). Observationally, the present-day O2s of rocky bodies in the solar system spans several orders of magnitude from roughly 5 units below IW (Mercury) to 2 units above IW (carbonaceous chondrites) (e.g., Grewal et al., 2024), with several extrasolar rocky bodies showing comparable redox states (Doyle et al., 2019). Early Earth conditions during the magma ocean stage were close to IW (Sossi et al., 2020).
@@
Therefore, to capture this diversity, we centre our log-uniform prior on zero offset relative to IW (anchored to early Earth) and allow sampling across a broad range of units. While no prior is without limitations, we adopt this scale-invariant formulation because it more faithfully represents ignorance over the parameter’s order of magnitude. It offers the dual advantages of anchoring to a well-motivated reference point (early Earth) while permitting exploration beyond the present-day solar system constraints on O2, which is warranted by the potentially greater diversity of exoplanetary environments. For consistency with the log-uniform prior, we apply the same sampling strategy to the other two parameters: the total hydrogen inventory and the C/H mass ratio are each varied across one order of magnitude on either side of an Earth-like reference value. The upper bound of the total hydrogen inventory is motivated by planet formation simulations, which show that terrestrial planets can accrete several times Earth’s water content (Raymond et al., 2007). The lower bound, by contrast, reflects scenarios of limited delivery or substantial early loss.
@@
The corner plots display bivariate scatter plots (e.g., Figure 1), which directly reflect the model outcomes and are therefore insensitive to the choice of prior, depending only on the sampled parameter ranges. In contrast, the likelihood contours, marginal distributions, and derived summary statistics (e.g., Table 3) are prior-dependent. A uniform distribution might appear to be a simple alternative prior, but its inadequacy for O2 is evident: when sampled uniformly between and , the mean value in space is , heavily skewed toward the upper bound and far outside the range characteristic of the early Earth or any known rocky body in the solar system. New observations and models may eventually reveal the true distributions of the parameters that govern atmospheric diversity, but in the absence of such knowledge, a logarithmic sampling strategy remains the most reasonable and defensible choice.
@@ Appendix B Partially molten mantle
We repeat the analysis presented in Section 3.1 for a partially molten TRAPPIST-1e with a mantle melt fraction of 10%, whereby the influence of solubility is lessened compared to a fully molten mantle (compare Figure 1, 2, 4, 5 to Figure B1, B2, B3, B4, respectively). Even with just 10% melt fraction, C/H in the atmosphere is boosted and its distribution broadened compared to total C/H (Figure 3). Compared to a fully molten mantle, a partially molten mantle permits a greater diversity of atmospheres since a broader range of gas mixtures comprised of hydrogen and carbon species are possible. Although larger H2O atmospheres can occur, the median H2O in the atmosphere is still only 11 mol.% versus 1 mol.% for a fully molten mantle. The mean molar mass (Figure B1, bottom right) reflects this greater atmospheric diversity, where a subsidiary peak around 5 g/mol represents H2-dominated atmospheres at reduced conditions in a mixture with CH4 and CO. The CO peak at 28 g/mol is evident, albeit now supplemented by mixtures of H2O (18 g/mol) and CO2 (44 g/mol) in approximately equal proportion. The molar mass distribution also has an extended tail, driven by the prevalence of CO across a broad range of oxygen fugacity as well as the emergence of CO2 (44 g/mol) and SO2 (64 g/mol) as prevalent oxidized species. Regardless, the thermodynamic stability of CO over a range of O2 ensures that it remains a dominant species.
@@
For a partially molten mantle, dissolved H is more than 60% for models above the IW buffer due to abundant H2O which is soluble. For reduced conditions below the IW buffer H2 is more abundant than H2O, and since H2 is not as soluble as H2O less total hydrogen is dissolved, between about 10% and 60%. For comparison, dissolved C never exceeds around 3% regardless of whether the reduced (CO) or oxidized (CO2) species is most abundant. We recall that total C/H is an input parameter, which is varied between -1 and 1 log10 units by mass, which corresponds approximately to between -2 and 0 log10 units by moles. This gives rise to a median log10 C/H in the atmosphere of -0.2 and C/H can reach several hundred for the most oxidized conditions. Hence even with a small mantle melt fraction (10%) total C/H can differ from atmospheric C/H by around two orders of magnitude.
@@
Figure B4 shows the atmospheres at 280 K derived from the partially molten starting state. Compared to those derived from the fully molten starting state, the atmospheres exhibit the same complex relationships albeit the solution space is enlarged for certain quantities. This is because the smaller initial C/H allows H2 and CH4 atmospheres to be more prevalent at reduced starting conditions which can produce mixtures of hydrogen and carbon species. Nevertheless, CO2 atmospheres are dominant in similarity with atmospheres derived from the fully molten starting state. H2O is only present in the atmosphere below the percent level since water condenses to produce a median surficial reservoir of 0.4 Earth oceans by mass (EO) with a maximum inventory around 5.5 EO, which is almost a ten fold increase compared to the fully molten starting condition (Figure 5). -sulfur condensation can produce up to around 0.4 EO, although an expected amount is closer to 0.1 EO. In contrast, the total graphite reservoir is usually around 0.3 EO, but for reduced starting conditions can reach up to 10 EO. Large reservoirs of all three condensates can coexist, usually for starting conditions around the IW buffer.
@@
The overall trends of atmospheres above planetary surfaces with stable condensates are comparable to the fully molten starting state in terms of the dominance of CO2 and CH4 atmospheres (compare Figure B5 and 9). However, there are a few notable differences. First, most water oceans are stable beneath CO2 atmospheres and most graphite reservoirs are stable beneath CH4 atmospheres. This trend is reversed compared to the atmospheres derived from a fully molten starting state. Second, H2S atmospheres are absent for the fully molten starting state, but now feature above all three condensate reservoirs. Third, SO2 atmospheres can support up to 3 EO of water as well as 0.15 EO of -sulfur. Finally, N2 atmospheres, although low in number, are also compatible with large reservoirs of water and -sulfur. Figure B6 shows the atmospheric speciation in equilibrium with stable reservoirs of water and graphite only, which excludes -sulfur and ammonium chloride compared to Figure 10 .
@@
Tables B1 and B2 summarise the atmosphere speciation for TRAPPIST-1e at 280 K for atmospheres with stable water, graphite, and -sulfur, for fully molten and partially molten starting states, respectively (Figure 10).
@@
We independently verified the results of our model by comparison to FactSage calculations (version 8.2, Bale et al., 2016), assuming ideal gas behaviour (Table C1). This comparison reveals that the partial pressure of a given volatile differs by at most a few percent. Additional comparison cases are available in the open source code including cases with condensates.
@@ Appendix D Bounded real gas equation of state
Real gas equations of state (EOS) are usually calibrated with experimental data and therefore valid within a certain range of temperature and pressure. Even if the parameters of a real gas EOS are within the calibrated range, numerical solvers perform better when the function is both bounded and smooth because these properties help avoid instability and ensure reliable convergence. Boundedness ensures that the solution doesn’t experience unphysical behaviours, such as division by zero or infinite values, which can lead to numerical errors or divergence. Smoothness, on the other hand, helps prevent large, abrupt changes in the equation of state that can cause solvers to struggle with step-size control or lead to inaccurate approximations, ultimately improving the robustness and accuracy of the numerical methods.
For a given EOS, a pressure range is defined based on calibration data to bracket where the EOS is physically reasonable and well-behaved. Below the minimum calibration pressure, often around 1 bar, the gas species is set to obey the ideal gas law. Above the maximum calibration pressure , we let the compressibility factor (e.g., Equation 1) follow an empirical virial-like equation (e.g., Saxena & Fei, 1987a, b):
where is the gas constant, temperature, molar volume, the compressibility factor of the EOS evaluated at (, T), and the gradient of with respect to above , which could be ignored (set to zero) for simplicity. The volume integral between and can be expressed analytically:

### v3

@@ 5 Conclusions
For Earth-sized rocky planets such as TRAPPIST-1e, the atmosphere during the magma ocean stage is likely CO-rich due to two key factors. First, CO(g) is thermodynamically favored at 1800 K across a broad range of oxygen fugacities. Second, carbon-bearing species are significantly less soluble than hydrogen-bearing species—particularly H2O(g)—,leading to an elevated atmospheric C/H, often exceeding the bulk C/H by several orders of magnitude. As a result, hydrogen-bearing species are depleted in the atmosphere, even when the mantle is only partially molten. In a fully molten mantle, the further depletion of hydrogen by the sequestration of H2O(g) into the liquid leads to atmospheres that are defined by entirely continuous mixing ratios of their constituent species, favoring atmospheres dominated by (1) heavy carbon species (predominantly CO(g)), (2) CO2(g) SO2(g)-bearing varieties at high O2s (IW = +3 to +4), or (3) H2(g) CH4(g)-bearing types at low O2 (below IW = -3).
@@
Based on the equilibrium results of atmospheres with molten mantles, we cool the atmospheres to 280 K and permit (if thermodynamically stable) the formation of water (oceans), graphite, -sulfur, and ammonium chloride. Unlike at high temperatures, the presence of condensates buffers gas mixing ratios to produce discrete ”families” of atmospheres. Notably, CO2-rich atmospheres are not only widespread (83% of cases) but also uniquely capable of supporting a wider variety of surface condensates compared to those dominated by other species. A small fraction (5%) are CH4-rich atmospheres that can sustain both water and graphite, while roughly 8% are CO2–SO2-rich in equilibrium with native sulfur. The remainder (1%) are H2-rich. Depending on the conditions during early equilibration with a magma ocean, atmospheres rich in N2, HCl, or H2S can also host diverse surface condensates; however, such cases comprise only 1% of all cases. In short, water condensation buffers the partial pressures of hydrogen-bearing species in the atmosphere, such that even low C/H does not produce substantial hydrogen atmospheres except at the most reduced conditions where H2 can dominate. Thus, for both high and low temperature atmospheres, carbon always wins.
@@
Our simulations can be considered as geological reference cases—the expected state of a rocky planet after its earliest evolution. From this state, subsequent processes such as geological outgassing and atmospheric (photo)chemistry, dynamics, and escape can modify the atmosphere over geological timescales. Hence, our results can be used to inform the priors and initial conditions of other models, as well as provide an observational baseline for rocky planets to determine if they have evolved away from their birth state. In this regard, the simulations can inform observational studies and instrument requirements to investigate the atmospheres of temperate rocky planets.
@@
For sub-Neptune-sized planets, we performed nonideal gas-phase chemistry simulations including dominant C- and H-bearing species at conditions relevant to the magma ocean–envelope interface. The combination of solubility and nonideality conspire to stabilize CH4 as a major gas species for a range of parameter choices that are reasonable given current observational constraints for sub-Neptunes. As for Earth-sized rocky planets, this result is a natural outcome of the high solubility of hydrogen-bearing species (notably molecular hydrogen for sub-Neptunes) compared to carbon-bearing species, a result that is exacerbated due to nonideality at the high-temperature (3000 K) and pressure (several GPa) conditions assumed to occur at the interface. Furthermore, the transition between a H2 and CH4 atmosphere can occur within the range of observational uncertainties in mass and radius, even for a fixed volatile budget and O2, thereby providing a new challenge to infer atmospheric composition. Given detections of carbon-bearing species in the atmospheres of sub-Neptunes, such as K2-18b and TOI-270d, our findings emphasize the need to include and realistically model gas envelopes throughout their extent.
@@
To achieve these results, we devised an open-source code Atmodeller, built on JAX, to probe the nature of atmospheres above rocky planets from Earth size to sub-Neptune size. The code includes solubility laws, real gas EOS, and thermodynamic data. Leveraging the functionality of the JAX Python library allows the code to be both user friendly and highly performant, and community development of the code is encouraged.
acknowledgments
@@
We thank Daniel Kitzmann, Patrick Kidger, and Johanna Haffner for discussions. D.J.B., M.A.T., and P.A.S. were supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract No. MB22.00033, a SERI-funded ERC Starting grant “2ATMO.” P.A.S. also thanks the Swiss National Science Foundation (SNSF) through an Eccellenza Professorship (#203668). M.A.T. is supported by NASA through the NASA Hubble Fellowship grant #HST-HF2-51545 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. K.H. acknowledges the FED-tWIN research program STELLA (Prf-2021-022) funded by the Belgian Science Policy Office (BELSPO) and the FWO research grant G014425N. M.T. acknowledges financial support from the Chair of Theoretical Astrophysics of Extrasolar Planets at LMU Munich.
@@ Appendix A Sampling Strategy and Parameter Choices
The logarithmic sampling strategy adopted to explore the atmospheric diversity of TRAPPIST-1e is physically motivated as follows. Oxygen fugacity (O2) exerts a fundamental control on atmospheric composition, particularly through its influence on redox speciation; for example, determining the relative abundances of reduced versus oxidized species such as H2 and H2O. Thermodynamically, O2 is defined by an equilibrium reaction, and when referenced to Fe–FeO (IW) equilibrium (Equation E1) is conventionally expressed as a offset relative to the IW buffer. This convention naturally supports the use of a log-uniform prior, especially within a thermochemical framework where partial pressures, fugacities, and equilibrium constants combine multiplicatively (and thus additively in log-space). Observationally, the present-day O2s of rocky bodies in the solar system spans several orders of magnitude from roughly 5 units below IW (Mercury) to 2 units above IW (carbonaceous chondrites) (e.g., Grewal et al., 2024), with several extrasolar rocky bodies showing comparable redox states (Doyle et al., 2019). Early Earth conditions during the magma ocean stage were close to IW (Sossi et al., 2020).
@@
Therefore, to capture this diversity, we center our log-uniform prior on zero offset relative to IW (anchored to early Earth) and allow sampling across a broad range of units. While no prior is without limitations, we adopt this scale-invariant formulation because it more faithfully represents ignorance over the parameter’s order of magnitude. It offers the dual advantages of anchoring to a well-motivated reference point (early Earth) while permitting exploration beyond the present-day solar system constraints on O2, which is warranted by the potentially greater diversity of exoplanetary environments. For consistency with the log-uniform prior, we apply the same sampling strategy to the other two parameters: the total hydrogen inventory and the C/H mass ratio are each varied across one order of magnitude on either side of an Earth-like reference value. The upper bound of the total hydrogen inventory is motivated by planet formation simulations, which show that terrestrial planets can accrete several times Earth’s water content (Raymond et al., 2007). The lower bound, by contrast, reflects scenarios of limited delivery or substantial early loss.
@@
The corner plots display bivariate scatter plots (e.g., Figure 1), which directly reflect the model outcomes and are therefore insensitive to the choice of prior, depending only on the sampled parameter ranges. In contrast, the likelihood contours, marginal distributions, and derived summary statistics (e.g., Table 3) are prior dependent. A uniform distribution might appear to be a simple alternative prior, but its inadequacy for O2 is evident: when sampled uniformly between and , the mean value in space is 4.7, which is heavily skewed toward the upper bound and far outside the range characteristic of the early Earth or any known rocky body in the solar system. New observations and models may eventually reveal the true distributions of the parameters that govern atmospheric diversity, but, in the absence of such knowledge, a logarithmic sampling strategy remains the most reasonable and defensible choice.
@@ Appendix B Partially molten mantle
We repeat the analysis presented in Section 3.1 for a partially molten TRAPPIST-1e with a mantle melt fraction of 10%, whereby the influence of solubility is lessened compared to a fully molten mantle (compare Figures 1, 2, 4, 5 to Figures B1, B2, B3, B4, respectively). Even with just 10% melt fraction, C/H in the atmosphere is boosted and its distribution broadened compared to total C/H (Figure 3). Compared to a fully molten mantle, a partially molten mantle permits a greater diversity of atmospheres since a broader range of gas mixtures comprised of hydrogen and carbon species are possible. Although larger H2O atmospheres can occur, the median H2O in the atmosphere is still only 11 mol% versus 1 mol% for a fully molten mantle. The mean molar mass (Figure B1, bottom right) reflects this greater atmospheric diversity, where a subsidiary peak around 5 g mol-1 represents H2-dominated atmospheres at reduced conditions in a mixture with CH4 and CO. The CO peak at 28 g mol-1 is evident, albeit now supplemented by mixtures of H2O (18 g mol-1) and CO2 (44 g mol-1) in approximately equal proportion. The molar mass distribution also has an extended tail, driven by the prevalence of CO across a broad range of oxygen fugacity as well as the emergence of CO2 (44 g mol-1) and SO2 (64 g mol-1) as prevalent oxidized species. Regardless, the thermodynamic stability of CO over a range of O2 ensures that it remains a dominant species.
@@
For a partially molten mantle, dissolved H is more than 60% for models above the IW buffer due to abundant H2O which is soluble. For reduced conditions below the IW buffer, H2 is more abundant than H2O, and, since H2 is not as soluble as H2O, less total hydrogen is dissolved, between about 10% and 60%. For comparison, dissolved C never exceeds around 3% regardless of whether the reduced (CO) or oxidized (CO2) species is most abundant. We recall that total C/H is an input parameter, which is varied between -1 and 1 log10 units by mass, which corresponds approximately to between -2 and 0 log10 units by moles. This gives rise to a median log10 C/H in the atmosphere of -0.2 and C/H can reach several hundred for the most oxidized conditions. Hence even with a small mantle melt fraction (10%) total C/H can differ from atmospheric C/H by around two orders of magnitude.
@@
Figure B4 shows the atmospheres at 280 K derived from the partially molten starting state. Compared to those derived from the fully molten starting state, the atmospheres exhibit the same complex relationships, albeit the solution space is enlarged for certain quantities. This is because the smaller initial C/H allows H2 and CH4 atmospheres to be more prevalent at reduced starting conditions which can produce mixtures of hydrogen and carbon species. Nevertheless, CO2 atmospheres are dominant in similarity with atmospheres derived from the fully molten starting state. H2O is only present in the atmosphere below the percent level since water condenses to produce a median surficial reservoir of 0.4 Earth oceans by mass (EO) with a maximum inventory around 5.5 EO, which is almost a tenfold increase compared to the fully molten starting condition (Figure 5). -sulfur condensation can produce up to around 0.4 EO, although an expected amount is closer to 0.1 EO. In contrast, the total graphite reservoir is usually around 0.3 EO, but for reduced starting conditions can reach up to 10 EO. Large reservoirs of all three condensates can coexist, usually for starting conditions around the IW buffer.
@@
The overall trends of atmospheres above planetary surfaces with stable condensates are comparable to the fully molten starting state in terms of the dominance of CO2 and CH4 atmospheres (compare Figures B5 and 9). However, there are a few notable differences. First, most water oceans are stable beneath CO2 atmospheres and most graphite reservoirs are stable beneath CH4 atmospheres. This trend is reversed compared to the atmospheres derived from a fully molten starting state. Second, H2S atmospheres are absent for the fully molten starting state, but now feature above all three condensate reservoirs. Third, SO2 atmospheres can support up to 3 EO of water as well as 0.15 EO of -sulfur. Finally, N2 atmospheres, although low in number, are also compatible with large reservoirs of water and -sulfur. Figure B6 shows the atmospheric speciation in equilibrium with stable reservoirs of water and graphite only, which excludes -sulfur and ammonium chloride compared to Figure 10 .
@@
Tables B1 and B2 summarize the atmosphere speciation for TRAPPIST-1e at 280 K for atmospheres with stable water, graphite, and -sulfur, for fully molten and partially molten starting states, respectively (Figure 10).
@@
We independently verified the results of our model by comparison to FactSage calculations (version 8.2, Bale et al., 2016), assuming ideal gas behavior (Table C1). This comparison reveals that the partial pressure of a given volatile differs by at most a few percent. Additional comparison cases are available in the open-source code including cases with condensates.
@@ Appendix D Bounded real gas equation of state
Real gas EOS are usually calibrated with experimental data, and are therefore valid within a certain range of temperature and pressure. Even if the parameters of a real gas EOS are within the calibrated range, numerical solvers perform better when the function is both bounded and smooth because these properties help avoid instability and ensure reliable convergence. Boundedness ensures that the solution does not experience unphysical behaviors, such as division by zero or infinite values, which can lead to numerical errors or divergence. Smoothness, on the other hand, helps prevent large, abrupt changes in the EOS that can cause solvers to struggle with step-size control or lead to inaccurate approximations, ultimately improving the robustness and accuracy of the numerical methods.
For a given EOS, a pressure range is defined based on calibration data to bracket where the EOS is physically reasonable and well behaved. Below the minimum calibration pressure, often around 1 bar, the gas species is set to obey the ideal gas law. Above the maximum calibration pressure , we let the compressibility factor (e.g., Equation 1) follow an empirical virial-like equation (e.g., Saxena & Fei, 1987a, b):
where is the gas constant, temperature, molar volume, the compressibility factor of the EOS evaluated at (, T), and the gradient of with respect to above , which could be ignored (set to zero) for simplicity. The volume integral between and can be expressed analytically